Calculus I, Fall 2012 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 05

1. Section 3.8: Read problem 24.

Also read page 190 about the limitation of Newton’s method. (i.e. When does it work and not work?)

2. Section 3.8: problems 11, 13, 22. Just write the formula of Newton’s iteration. Need not get the numerical values for problems 13, 22.

3. Section 4.2: problems 31, 33, 57, 61, 63, 81.